# Can Large Language Models be Good Path Planners? A Benchmark and Investigation on Spatial-temporal Reasoning

### v1

@@ B.3 Chain of Thought Prompt
{tcolorbox}
[enhanced,attach boxed title to top center=yshift=-3mm,yshifttext=-1mm, colback=black!5!white,colframe=black!75!black,colbacktitle=white!95!black, title=CoT Prompt,fonttitle=,coltitle=black, boxed title style=size=small,colframe=black!50!black ,before upper= ] Provide a sequence of actions to navigate a world to reach a goal similarly to the examples below. (0,0) is located in the upper-left corner and (M, N) lies in the M row and N column. ###
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,1). Go from (0,1) to (3,4)
Actions: (3,4) is 3 steps down and 3 steps to the right of (0,1). To avoid the obstacle at (2,1), which is 2 steps down from (0,1), I should start by moving right. Therefore, my action sequence is: right right right down down down.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,5) and (1,2). Go from (5,4) to (0,5)
Actions: (0,5) is 5 steps up and 1 step to the right of (5,4). To avoid the obstacle at (1,5), which is 4 steps up and 1 step to the right from (5,4), I should move right last. Therefore, my action sequence is: up up up up up right.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,3), (2,5) and (5,2). Go from (4,2) to (0,5)
Actions: (0,5) is 4 steps up and 3 steps to the right of (4,2). I can start by going up. To avoid the obstacle at (0,3), which is 4 steps up and 1 step to the right of (4,2), I should take a right at (1,3). Therefore my action sequence is: up up up right right up right
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (3,5), (4,2), (3,3) and (0,4). Go from (1,5) to (3,1)
Actions: (3,1) is 4 steps to the left and 2 steps to the right of (1,5). No obstacles fall on this path. Therefore my action sequence is: left left left left down down.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,5), (5,2), (0,4), (1,4) and (0,1). Go from (4,2) to (1,2)
Actions: (1,2) is 3 steps up from (4,2). No obstacles fall on this path. Therefore my action sequence is: up up up.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,4) and (1,5). Go from (0,5) to (1,1)
Actions: (0,5) is surrounded by obstacles. Therefore, the goal is not reachable from my location.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,4), (4,4), (5,3) and (4,5). Go from (0,4) to (5,5)
Actions: (5,5) is surrounded by obstacles. Therefore, the goal is not reachable from my location.
@@ B.4 ReAct Prompt
{tcolorbox}
[enhanced,attach boxed title to top center=yshift=-3mm,yshifttext=-1mm, colback=black!5!white,colframe=black!75!black,colbacktitle=white!95!black, title=ReAct Prompt,fonttitle=,coltitle=black, boxed title style=size=small,colframe=black!50!black, before upper= ]
Provide a sequence of actions to navigate a world to reach a goal similarly to the examples below. (0,0) is located in the upper-left corner and (M, N) lies in the M row and N column.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,3), (5,5) and (5,2). Go from (0,4) to (5,0)
Thought 1: (5,0) is 5 steps down and 4 steps to the left of (0,4). To avoid the obstacle at (2,3), which is 2 steps down and 1 step to the left from (0,4), I should move left first.
Act 1: left left left left down down down down down.
Obs 1: Performing the action sequence leads to (5,0). The task has been solved.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (3,0), (1,5), (3,4), (5,2) and (5,3). Go from (1,4) to (4,0)
Thought 1: (4,0) is 3 steps down and 4 steps to the left of (1,4). To avoid the obstacle at (3,0), which is 2 steps down and 4 steps to the left from (1,4), I should move left last.
Act 1: down down down left left left left.
Obs 1: After executing the first step, I am at (2,4). If I execute the next step I will run into the obstacle at (4,3).
Thought 2: I have to find a path to get to (4,0) from (2,4). (4,0) is 2 steps down and 4 steps to the left from (2,4). In order to avoid the obstacle at (4,3), which is one step down, I have to start by moving left first.
Act 2: left left left left down down
Obs 2: After executing the first 4 steps, I am at (2,0). If I execute the next step I will run into the obstacle at (3,0).
Thought 3: I have to find a path to get to (4,0) from (2,0). (4,0) is 2 steps down from (2,0). In order to avoid the obstacle at (4,3), which is one step down, I have to move right, then take two steps down, then move left.
Act 3: right down down left
Obs 3: Performing the action sequence leads to (4,0). The task has been solved.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,5), (3,5), (0,0), (4,5) and (4,4). Go from (1,1) to (5,0)
Thought 1: (5,0) is 4 steps down and 1 step to the left of (1,1). To avoid the obstacle at (0,0), which is 1 step up and 1 step to the left from (1,1), I should move down first.
Act 1: down down down down left.
Obs 1: Performing the action sequence leads to (5,0). The task has been solved.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (5,5), (5,0), (3,5) and (4,0). Go from (3,4) to (1,3)
Thought 1: (1,3) is 2 steps up and 1 step to the left of (3,4). No obstacles fall on this path.
Act 1: up up left
Obs 1: Performing the action sequence leads to (1,3). The task has been solved.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,3), (4,2), (3,3) and (1,4). Go from (3,2) to (4,4)
Thought 1: (4,4) is 1 step down and 2 step to the right of (3,2). To avoid the obstacle at (3,3) which is one step to the right, I have to go down first
Act 1: down left left
Obs 1: If I execute the first step I will run into the obstacle at (4,2).
Thought 2: I have to find a path to get to (4,4) from (3,2). (4,4) is 1 step down and 2 step to the right of (3,2). In order to avoid the obstacle at (4,2) which is one step down, and the obstacle at (3,3) which one step to the right, I have to move up, take two steps to the right and two steps down.
Act 2: up right right down down
Obs 2: Performing the action sequence leads to (4,4). The task has been solved.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,3), (1,2), (3,5) and (0,1). Go from (0,2) to (3,1)
Thought 1: (3,1) is 3 steps down and 1 step to the left of (0,2). To avoid the obstacle at (1,2), which is 1 step down from (0,2), I should start by moving down.
Act 1: down down down left
Obs 1: If I execute the first step I will run into the obstacle at (1,2).
Thought 2: (0,2) is surrounded by obstacles. Therefore, the goal is not reachable from my location.
Act 2: No action
Obs 2: No action is to be performed. The goal is not reachable. The task has been solved.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,5), (5,2), (0,4), (1,4) and (0,1). Go from (4,1) to (1,5)
Thought 1: (1,5) is 3 steps up and 4 steps to the right of (4,1). To avoid the obstacle at (2,5), which is 2 steps up and 4 steps to the right from (4,1), I should move right last.
Act 1: up up up right right right right
Obs 1: After executing the first 5 steps, I am at (1,3). If I execute the next step I will run into the obstacle at (1,4).
Thought 2: (1,5) is surrounded by obstacles. Therefore, the goal is not reachable from my location.
Act 2: No action
Obs 2: No action is to be performed. The goal is not reachable. The task has been solved.
@@ B.5 Ordering Prompts
{tcolorbox}
[enhanced,attach boxed title to top center=yshift=-3mm,yshifttext=-1mm, colback=black!5!white,colframe=black!75!black,colbacktitle=white!95!black, title=Prompting for a valid ordering,fonttitle=,coltitle=black, boxed title style=size=small,colframe=black!50!black, before upper= ]
@@
Provide a plan to navigate a world to reach all the goals while satisfying any constraints similarly to the examples below. (0,0) is located in the upper-left corner and (M, N) lies in the M row and N column.
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (5,2), (2,3) and (5,0). You are at (0,2). You have to visit p0, p1, p2, p3 and p4. p0 is located at (3,5), p1 is located at (5,4), p2 is located at (2,4), p3 is located at (3,2) and p4 is located at (4,4). Visit p1 and p3 before p0, p2 and p4.
Order: p3, p1, p4, p2, p0
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,5), (0,2) and (4,5). You are at (4,2). You have to visit p0, p1, p2, p3 and p4. p0 is located at (0,1), p1 is located at (2,2), p2 is located at (1,2), p3 is located at (5,3) and p4 is located at (5,5).
Order: p3, p4, p1, p2, p0
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,0), (3,3) and (1,1). You are at (3,5). You have to visit p0, p1, p2, p3, p4 and p5. p0 is located at (1,3), p1 is located at (0,4), p2 is located at (4,0), p3 is located at (2,4), p4 is located at (5,0) and p5 is located at (5,4). Visit p4, p3 and p2 before p0, p1 and p5.
Order: p3, p2, p4, p5, p0, p1
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,1). You are at (4,0). You have to visit p0, p1, p2, p3 and p4. p0 is located at (3,2), p1 is located at (1,1), p2 is located at (2,2), p3 is located at (0,4) and p4 is located at (1,5). Visit p2 and p3 before p0, p1 and p4
Order: p2, p3, p4, p1, p0
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,1). You are at (5,3). You have to visit p0 and p1. p0 is located at (2,5) and p1 is located at (2,2). Visit p1 before p0
Order: p1, p0
{tcolorbox}
[enhanced,attach boxed title to top center=yshift=-3mm,yshifttext=-1mm, colback=black!5!white,colframe=black!75!black,colbacktitle=white!95!black, title=Prompting for an optimal ordering,fonttitle=,coltitle=black, boxed title style=size=small,colframe=black!50!black, before upper= ]
@@
Provide an optimal plan to navigate a world to reach all the goals while satisfying any constraints similarly to the examples below. (0,0) is located in the upper-left corner and (M, N) lies in the M row and N column. A path is optimal if it satisfies the constraints using the minimum number of actions
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (5,2), (2,3) and (5,0). You are at (0,2). You have to visit p0, p1, p2, p3 and p4. p0 is located at (3,5), p1 is located at (5,4), p2 is located at (2,4), p3 is located at (3,2) and p4 is located at (4,4). Visit p1 and p3 before p0, p2 and p4.
Order: The optimal plan is: p3, p1, p4, p2, p0
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,5), (0,2) and (4,5). You are at (4,2). You have to visit p0, p1, p2, p3 and p4. p0 is located at (0,1), p1 is located at (2,2), p2 is located at (1,2), p3 is located at (5,3) and p4 is located at (5,5).
Order: The optimal plan is: p3, p4, p1, p2, p0
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,0), (3,3) and (1,1). You are at (3,5). You have to visit p0, p1, p2, p3, p4 and p5. p0 is located at (1,3), p1 is located at (0,4), p2 is located at (4,0), p3 is located at (2,4), p4 is located at (5,0) and p5 is located at (5,4). Visit p4, p3 and p2 before p0, p1 and p5.
Order: The optimal plan is: p3, p2, p4, p5, p0, p1
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,1). You are at (4,0). You have to visit p0, p1, p2, p3 and p4. p0 is located at (3,2), p1 is located at (1,1), p2 is located at (2,2), p3 is located at (0,4) and p4 is located at (1,5). Visit p2 and p3 before p0, p1 and p4
Order: The optimal plan is: p2, p3, p4, p1, p0
Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,1). You are at (5,3). You have to visit p0 and p1. p0 is located at (2,5) and p1 is located at (2,2). Visit p1 before p0
Order: The optimal plan is: p1, p0
@@ Appendix D Additional Analysis
Are the fine-tuned models better at generalizing to unseen placements? We also look at whether the fine-tuned models are better at generalizing to novel initial and goal placements within environments used for training as opposed to generalizing to unseen environments. Our results, presented in 11, show that the models achieve almost the same performance on this test; indicating that these models do not overfit to the environments they are trained on.

### v2

@@ B.3 Chain of Thought Prompt
CoT Prompt Provide a sequence of actions to navigate a world to reach a goal similarly to the examples below. (0,0) is located in the upper-left corner and (M, N) lies in the M row and N column. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,1). Go from (0,1) to (3,4) Actions: (3,4) is 3 steps down and 3 steps to the right of (0,1). To avoid the obstacle at (2,1), which is 2 steps down from (0,1), I should start by moving right. Therefore, my action sequence is: right right right down down down. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,5) and (1,2). Go from (5,4) to (0,5) Actions: (0,5) is 5 steps up and 1 step to the right of (5,4). To avoid the obstacle at (1,5), which is 4 steps up and 1 step to the right from (5,4), I should move right last. Therefore, my action sequence is: up up up up up right. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,3), (2,5) and (5,2). Go from (4,2) to (0,5) Actions: (0,5) is 4 steps up and 3 steps to the right of (4,2). I can start by going up. To avoid the obstacle at (0,3), which is 4 steps up and 1 step to the right of (4,2), I should take a right at (1,3). Therefore my action sequence is: up up up right right up right ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (3,5), (4,2), (3,3) and (0,4). Go from (1,5) to (3,1) Actions: (3,1) is 4 steps to the left and 2 steps to the right of (1,5). No obstacles fall on this path. Therefore my action sequence is: left left left left down down. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,5), (5,2), (0,4), (1,4) and (0,1). Go from (4,2) to (1,2) Actions: (1,2) is 3 steps up from (4,2). No obstacles fall on this path. Therefore my action sequence is: up up up. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,4) and (1,5). Go from (0,5) to (1,1) Actions: (0,5) is surrounded by obstacles. Therefore, the goal is not reachable from my location. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,4), (4,4), (5,3) and (4,5). Go from (0,4) to (5,5) Actions: (5,5) is surrounded by obstacles. Therefore, the goal is not reachable from my location.
@@ B.4 ReAct Prompt
ReAct Prompt Provide a sequence of actions to navigate a world to reach a goal similarly to the examples below. (0,0) is located in the upper-left corner and (M, N) lies in the M row and N column. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,3), (5,5) and (5,2). Go from (0,4) to (5,0) Thought 1: (5,0) is 5 steps down and 4 steps to the left of (0,4). To avoid the obstacle at (2,3), which is 2 steps down and 1 step to the left from (0,4), I should move left first. Act 1: left left left left down down down down down. Obs 1: Performing the action sequence leads to (5,0). The task has been solved. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (3,0), (1,5), (3,4), (5,2) and (5,3). Go from (1,4) to (4,0) Thought 1: (4,0) is 3 steps down and 4 steps to the left of (1,4). To avoid the obstacle at (3,0), which is 2 steps down and 4 steps to the left from (1,4), I should move left last. Act 1: down down down left left left left. Obs 1: After executing the first step, I am at (2,4). If I execute the next step I will run into the obstacle at (4,3). Thought 2: I have to find a path to get to (4,0) from (2,4). (4,0) is 2 steps down and 4 steps to the left from (2,4). In order to avoid the obstacle at (4,3), which is one step down, I have to start by moving left first. Act 2: left left left left down down Obs 2: After executing the first 4 steps, I am at (2,0). If I execute the next step I will run into the obstacle at (3,0). Thought 3: I have to find a path to get to (4,0) from (2,0). (4,0) is 2 steps down from (2,0). In order to avoid the obstacle at (4,3), which is one step down, I have to move right, then take two steps down, then move left. Act 3: right down down left Obs 3: Performing the action sequence leads to (4,0). The task has been solved. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,5), (3,5), (0,0), (4,5) and (4,4). Go from (1,1) to (5,0) Thought 1: (5,0) is 4 steps down and 1 step to the left of (1,1). To avoid the obstacle at (0,0), which is 1 step up and 1 step to the left from (1,1), I should move down first. Act 1: down down down down left. Obs 1: Performing the action sequence leads to (5,0). The task has been solved. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (5,5), (5,0), (3,5) and (4,0). Go from (3,4) to (1,3) Thought 1: (1,3) is 2 steps up and 1 step to the left of (3,4). No obstacles fall on this path. Act 1: up up left Obs 1: Performing the action sequence leads to (1,3). The task has been solved. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,3), (4,2), (3,3) and (1,4). Go from (3,2) to (4,4) Thought 1: (4,4) is 1 step down and 2 step to the right of (3,2). To avoid the obstacle at (3,3) which is one step to the right, I have to go down first Act 1: down left left Obs 1: If I execute the first step I will run into the obstacle at (4,2). Thought 2: I have to find a path to get to (4,4) from (3,2). (4,4) is 1 step down and 2 step to the right of (3,2). In order to avoid the obstacle at (4,2) which is one step down, and the obstacle at (3,3) which one step to the right, I have to move up, take two steps to the right and two steps down. Act 2: up right right down down Obs 2: Performing the action sequence leads to (4,4). The task has been solved. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,3), (1,2), (3,5) and (0,1). Go from (0,2) to (3,1) Thought 1: (3,1) is 3 steps down and 1 step to the left of (0,2). To avoid the obstacle at (1,2), which is 1 step down from (0,2), I should start by moving down. Act 1: down down down left Obs 1: If I execute the first step I will run into the obstacle at (1,2). Thought 2: (0,2) is surrounded by obstacles. Therefore, the goal is not reachable from my location. Act 2: No action Obs 2: No action is to be performed. The goal is not reachable. The task has been solved. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,5), (5,2), (0,4), (1,4) and (0,1). Go from (4,1) to (1,5) Thought 1: (1,5) is 3 steps up and 4 steps to the right of (4,1). To avoid the obstacle at (2,5), which is 2 steps up and 4 steps to the right from (4,1), I should move right last. Act 1: up up up right right right right Obs 1: After executing the first 5 steps, I am at (1,3). If I execute the next step I will run into the obstacle at (1,4). Thought 2: (1,5) is surrounded by obstacles. Therefore, the goal is not reachable from my location. Act 2: No action Obs 2: No action is to be performed. The goal is not reachable. The task has been solved.
@@ B.5 Ordering Prompts
Prompting for a valid ordering Provide a plan to navigate a world to reach all the goals while satisfying any constraints similarly to the examples below. (0,0) is located in the upper-left corner and (M, N) lies in the M row and N column. ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (5,2), (2,3) and (5,0). You are at (0,2). You have to visit p0, p1, p2, p3 and p4. p0 is located at (3,5), p1 is located at (5,4), p2 is located at (2,4), p3 is located at (3,2) and p4 is located at (4,4). Visit p1 and p3 before p0, p2 and p4. Order: p3, p1, p4, p2, p0 ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,5), (0,2) and (4,5). You are at (4,2). You have to visit p0, p1, p2, p3 and p4. p0 is located at (0,1), p1 is located at (2,2), p2 is located at (1,2), p3 is located at (5,3) and p4 is located at (5,5). Order: p3, p4, p1, p2, p0 ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,0), (3,3) and (1,1). You are at (3,5). You have to visit p0, p1, p2, p3, p4 and p5. p0 is located at (1,3), p1 is located at (0,4), p2 is located at (4,0), p3 is located at (2,4), p4 is located at (5,0) and p5 is located at (5,4). Visit p4, p3 and p2 before p0, p1 and p5. Order: p3, p2, p4, p5, p0, p1 ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,1). You are at (4,0). You have to visit p0, p1, p2, p3 and p4. p0 is located at (3,2), p1 is located at (1,1), p2 is located at (2,2), p3 is located at (0,4) and p4 is located at (1,5). Visit p2 and p3 before p0, p1 and p4 Order: p2, p3, p4, p1, p0 ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,1). You are at (5,3). You have to visit p0 and p1. p0 is located at (2,5) and p1 is located at (2,2). Visit p1 before p0 Order: p1, p0
@@
Prompting for an optimal ordering Provide an optimal plan to navigate a world to reach all the goals while satisfying any constraints similarly to the examples below. (0,0) is located in the upper-left corner and (M, N) lies in the M row and N column. A path is optimal if it satisfies the constraints using the minimum number of actions ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (5,2), (2,3) and (5,0). You are at (0,2). You have to visit p0, p1, p2, p3 and p4. p0 is located at (3,5), p1 is located at (5,4), p2 is located at (2,4), p3 is located at (3,2) and p4 is located at (4,4). Visit p1 and p3 before p0, p2 and p4. Order: The optimal plan is: p3, p1, p4, p2, p0 ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,5), (0,2) and (4,5). You are at (4,2). You have to visit p0, p1, p2, p3 and p4. p0 is located at (0,1), p1 is located at (2,2), p2 is located at (1,2), p3 is located at (5,3) and p4 is located at (5,5). Order: The optimal plan is: p3, p4, p1, p2, p0 ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (1,0), (3,3) and (1,1). You are at (3,5). You have to visit p0, p1, p2, p3, p4 and p5. p0 is located at (1,3), p1 is located at (0,4), p2 is located at (4,0), p3 is located at (2,4), p4 is located at (5,0) and p5 is located at (5,4). Visit p4, p3 and p2 before p0, p1 and p5. Order: The optimal plan is: p3, p2, p4, p5, p0, p1 ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (0,1). You are at (4,0). You have to visit p0, p1, p2, p3 and p4. p0 is located at (3,2), p1 is located at (1,1), p2 is located at (2,2), p3 is located at (0,4) and p4 is located at (1,5). Visit p2 and p3 before p0, p1 and p4 Order: The optimal plan is: p2, p3, p4, p1, p0 ### Task: You are in a 6 by 6 world. There are obstacles that you have to avoid at: (2,1). You are at (5,3). You have to visit p0 and p1. p0 is located at (2,5) and p1 is located at (2,2). Visit p1 before p0 Order: The optimal plan is: p1, p0
@@ Appendix D Additional Analysis
Are the fine-tuned models better at generalizing to unseen placements? We also look at whether the fine-tuned models are better at generalizing to novel initial and goal placements within environments used for training as opposed to generalizing to unseen environments. Our results, presented in 12, show that the models achieve almost the same performance on this test; indicating that these models do not overfit to the environments they are trained on.
@@
Can the results scale up to more realistic environments? Finally we look at whether the results we observed can be generalized to a more complex environment, that more closely resembles a real-world navigation scenario. We prompt GPT4-V(ision) using Naive Prompting, Action Effects Prompting and Chain-of-Thought Reasoning. The current version of GPT-4V struggled to learn from few-shot image-and-text examples, even when we tried providing the information over multiple conversational turns. As a result, all of our experiments were done in a zero-shot setting only. However, our findings from small 2D grid environments still held true here - we showed that (a) vanilla zero-shot GPT-4V just generates random action sequences, (b) augmenting GPT-4V with Action-Effect knowledge dramatically improves success rate, and (c) zero-shot GPT-4V with CoT gives similar output to GPT-4V Action-Effect due to the zero-shot limitation, but it also outlines a strategy of asking for environment feedback, further supporting the potential of ReAct.
Image: [Uncaptioned image] Naive Prompting: You are navigating the environment below, each elevated object is an obstacle that you have to avoid, Go from point 1 to point 2. Point 1 is located in the top left corner and point 2 is located in the bottom right corner. Your only available actions are (left, right, up, down). Your answer should consist of the predicted sequence solely. Naive Prompting Output: down, down, down, down, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left, left, down, down, right, right, down, down, left Remarks: Completely random sequence. Consistent with our finding that naive prompting fails at conceptualizing the task and environment.
Image: [Uncaptioned image] Action and Effects Prompting: You are navigating the environment below, each elevated object is an obstacle that you have to avoid, Go from point 1 to point 2. Your only available actions are (left, right, up, down). Find a path to go from point 1 to point 2. Point 1 is located in the top left corner and point 2 is located in the bottom right corner. Show me the predicted action sequence and what you think would happen with every action Action and Effects Prompting Output: Action: Move down. Outcome: You would move closer to the first obstacle row, avoiding collision. Action: Move down. Outcome: You would be positioned just before the first horizontal obstacle, still avoiding collision. Action: Move right. Outcome: You would navigate towards the gap between the first and second obstacle in the first row. Action: Move down. Outcome: You would now be aligned with the gap between the obstacles, moving closer to the second row. Action: Move right. Outcome: You would enter the second row, moving through the gap between the first and second obstacles. Action: Move down. Outcome: You would be positioned just before the obstacles in the second row. Action: Move right, and continue this action. Outcome: You would pass through the gaps between the obstacles in the subsequent rows, moving down each row closer to point 2. Action: Once aligned with point 2, move right. Outcome: You would reach point 2 located at the bottom right corner of the environment. Remarks: With this prompting method, GPT-4v is able to better reason about the environment and how to avoid obstacles, but the plan generated falls short, as it does not reach the goal.
Image: [Uncaptioned image] Chain of Thought Prompting: You are navigating the environment below, each elevated object is an obstacle that you have to avoid, Go from point 1 to point 2. Point 1 is located in the top left corner and point 2 is located in the bottom right corner. Your only available actions are (left, right, up, down). Provide a sequence of actions to reach point 2 from point 1. Think step-by-step Chain of Thought Prompting Output: Start at Point 1. Move right until the path is obstructed by an obstacle. When an obstacle is directly to the right, move down until you are able to move right again. Continue moving right until you encounter another obstacle. Repeat the process of moving down and then to the right until you can no longer move right because the path is either clear or because you have reached the last column before Point 2. Once in the column adjacent to Point 2, move down until you are aligned horizontally with Point 2. Move right into Point 2. Remarks: The predicted approach outlines a strategy based on environment feedback (e.g. Move right until the path is obstructed by an obstacle.), decomposing the task into segments that require moving in a single direction until an obstacle is encountered; which highlights that when these methods are augmented with a method similar to ReAct to identify when an obstacle has been encountered, it can successfully solve the task. Nevertheless, the solution generated is not optimal as it requires naively reasoning locally to avoid obstacles.